1) Supply and demand

P = 0.5QS + 30
P = -0.4QD + 120

a) Given the above equations, produce a chart illustrating both the supply and demand schedules in increments of 5 ranging from price = 50 to price = 110.
b) Solve for the equilibrium price and quantity and show your work.
c) Graph the result, labeling the axes, the supply and demand curves, the equilibrium point, and the price and quantity amounts. Use a proper scale.

d) Redo parts (a) through (c), this time reflecting a change in the supply curve to P = 0.5QS + 7.5.
e) Give two possible explanations for what could have caused such a change.
f) Now take the new graph (from part d) and assume that the government institutes a price floor equal to P = 80. Describe in detail what would be the impact on this market.

3) Welfare analysis
a) Explain in your own words the economic meaning of consumer surplus and

producer surplus
b) Calculate the consumer and producer surpluses from parts (a) through (c) of

question 1 and show on a separate graph.
c) Now explain clearly (and show all calculations) what happens to consumer surplus

and producer surplus as a result of the change depicted in question 1, part (d).

Show graphically.
d) Is there a deadweight loss? Why or why not? (If so, calculate it). Does society gain

or lose overall? Explain why this makes sense.
e) Compared to the situation in parts (c) and (d) of this question, what are the gains or

losses to consumers and producers from the price floor of 80. Show calculations. f) Is there a deadweight loss? Why or why not? And if so, calculate it.
g) Can we say definitively whether society overall benefited or lost? Explain.
Answer #1

As per the HOMEWORKLIB POLICY, answering the first four sub-parts

Page 1 SS : P-0.50+ 30 > QS= es >> OS = 2p60 DO : p= -0.400 +120 0.402= 120-P -> @d- 300 – 2.5p os od 40 175 60 70 55 162.5 1

d) Od = 300 - 2.5p Page 2 ss; P = 0,50$+7.5 = 10.50$ = P-7.5 >10% = 99-15 at equilibrium, ad = New as 5300 -2.50 = 22-15 => 4